The area of a triangle is 54 inches squared. Find the length of its sides if they form an arithmetic sequence.
Answer by Alan3354(69443)

The area of a triangle is 54 inches squared. Find the length of its sides if they form an arithmetic sequence.
==============
Area = b*h/2 = 54
b*h = 108
108 = 9*12
---> 9, 12 & 15